OpenStudy (anonymous):

translate "three plus two thirds ofa number is 7" into an equation and solve showing all work

OpenStudy (anonymous):

\[3+\frac{2}{3}x = 7\]

OpenStudy (anonymous):

3+2a/3=7 9+2a=21 2a=12 a=6